OpenStudy (anonymous):

but do the 2 stay or do they get delete out

OpenStudy (anonymous):

what I font get ur question

OpenStudy (anonymous):

*don't

OpenStudy (anonymous):

so the answer is X2 + 452=902

OpenStudy (anonymous):

x2+452=902 -452 -452 x2=450 divide and get 225

OpenStudy (anonymous):

\[what i was trying \to say is this X ^2+45^2=90^2 and that is the answer \right\]

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so i was right or wrong on the answer

OpenStudy (anonymous):

right